Given: You wish to have S60,000 to invest in 8 years at an annual interest of 7%. Find: The annual savings which needs to be deposited at the end of each year. Given: You wish to borrow $500,000 and the bank charges 15% interest monthly.
Answer #1

2.3

Annual interest rate R = 7%

Time n = 8 years

Future value required in 8 years = $60000

Let, annual savings = P

Then,

60000 = P*(1-1/1.07^8)/.07

60000 = P*5.9713

P = 60000/5.9713

P = $10048.06 or $10048

So, annual savings should be  $10048.06 or $10048.

Pl. repost other unanswered questions for their proper answers!